The nutritional formula Sustacal is supplied on 10-oz cans. How you prepare 40 oz of a half-strenght Sustacal solution
Answer #1

40 x 1/2 = 20 oz Sustacal.

Dulution is 40 - 20 = 20 oz.

20+20 = 40 oz.

So, take 20 oz of water and 2 cans of 10 oz of nutritional formula and add them. This will make half strength of 40 oz of a half-strenght Sustacal solution.